OpenStudy (anonymous):

simplify ((2+tan^2)/sec^2))-1. Please help

OpenStudy (anonymous):

\[\frac{2+\tan^2x}{\sec^2x}-1\\ \frac{1+1+\tan^2x}{\sec^2x}-1\\ \frac{1+\sec^2x}{\sec^2x}-1\\ \frac{1}{\sec^2x}+\frac{\sec^2x}{\sec^2x}-1\\ \cos^2x+1-1\\ \cos^2x \]